• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Meziani

    Log entry time 11:31:46 on July23,2005

    Entry number 149158

    keyword=Target status

    9:00 am Talked to CHL cryotech on call about the instability in
    CHL that was observed during the owl shift up to 8:30 am.
    Changes made to flow levels of ESR that were supposed to eleviate
    the problem.
    CHL flow seem stable since about 9:30 am